# Flexible inference of evolutionary accumulation dynamics using uncertain observational data

### v2

@@ Abstract
Understanding and predicting evolutionary accumulation pathways is a key objective in many fields of research, ranging from classical evolutionary biology to diverse applications in medicine. In this context, we are often confronted with the problem that data is sparse and uncertain. To use the available data as best as possible, inference approaches that can handle this uncertainty are required. One way that allows us to use not only cross-sectional data, but also phylogenetic related and longitudinal data, is using ‘hypercubic inference’ models. In this article we introduce HyperLAU, a new algorithm for hypercubic inference that makes it possible to use datasets including uncertainties for learning evolutionary pathways. Expanding the flexibility of accumulation modelling, HyperLAU allows us to infer dynamic pathways and interactions between features, even when large sets of particular features are unobserved across the source dataset. We show that HyperLAU is able to highlight the main pathways found by other tools, even when up to 50% of the features in the input data are uncertain. Additionally, we demonstrate how it can help to overcome possible biases that can occur then reducing the used data by excluding uncertain parts. We illustrate the approach with a case study on multidrug resistance in tuberculosis, showing that HyperLAU allows more flexible data and provides new information about evolutionary pathways compared to existing approaches.
@@ 1 Introduction
Many processes in the natural sciences and especially in biology and medicine can be modelled as the accumulation or loss of binary features [Diaz-Uriarte and Johnston, 2025, Schill et al., 2024b, Diaz-Uriarte and Herrera-Nieto, 2022]. Common examples include evolutionary processes, such as accumulation of mutations in cancer [Schwartz and Schäffer, 2017, Diaz-Uriarte and Vasallo, 2019, Beerenwinkel et al., 2015, Takeshima and Ushijima, 2019, Moen and Johnston, 2023], gene loss in mitochondria [Johnston and Williams, 2016, Maier et al., 2013], mutations in bacteria [Nichol et al., 2015, Tan et al., 2011, Greenbury et al., 2020], and the appearance of other phenotypic characters [Moen and Johnston, 2023, Nichol et al., 2015]. Other examples of accumulation processes include, for example, patients acquiring symptoms in progressing diseases [Dalgıç et al., 2021, Johnston et al., 2019, Schill et al., 2024a]. Understanding and predicting such processes can contribute both to the basic knowledge of the mechanisms involved, and to potential application – for example, in triaging patients [Johnston et al., 2019] and determining optimal treatments [Aga et al., 2024, Renz et al., 2024].
To address this problem, many different types of mathematical models for accumulation processes have been derived over the last decades, many of them using discrete binary structures. Among the first models in this context, the Mk (Markov k-state) model [Pagel, 1994, Lewis, 2001] for the evolution of traits across a group of related species were introduced. Recently, [Johnston and Diaz-Uriarte, 2024] demonstrated how the Mk model can be used to infer accumulation processes, including the possibility of reversible transitions.
@@
In several scientific and medical contexts, data on accumulation processes often includes uncertainties and unknown states. The extent to which existing models can handle this uncertainty is typically limited. In most of the named models, there are only general correction mechanisms or error rates included, which can account for some noise. MO [Gao et al., 2022] explicitly allows missing or ambiguous information for specific features, but is subject to the infinite-sites assumption (every event occurs at most once, prohibiting parallel evolution), which is not appropriate in several scientific contexts. HyperTraPS-CT [Aga et al., 2024] does not have this assumption, but represents data as a set of ancestor-descendant states and can only naturally deal with uncertainties in the descendant states, not in the ancestral ones. So far the only model that can fulfill all these requirements is the Mk-model [Johnston and Diaz-Uriarte, 2024] which, however, is limited to very few features due to the high computational runtime.
Here, we introduce HyperLAU (Hypercubic transition paths with Linear Algebra for Uncertainty), an approach for accumulation modelling that has no underlying infinite site assumption but can deal with longitudinal input data, including specified uncertainties in ancestor and descendant states. On the same time, it can handle input data with more features than the Mk-model in a smaller amount of time. Pursuing the idea of an hypercubic transition graph and maintaining the option of longitudinal and/or phylogenetic data as input, we developed an approach where precise positions of uncertainty can be specified in both the ancestor and descendant strings. In the center of the algorithm are simple matrix-vector multiplications, as known from linear algebra, which makes the approach intuitive and easy understandable (see Figure 1). Borrowing the idea for the option to model different orders of dependencies from [Greenbury et al., 2020], we additionally keep this flexibility, what is a further advantage of our model.
We will particularly motivate the approach in the context of anti-microbial resistance (AMR) evolution – the evolution of pathogens to different drugs. AMR is one of the top threats to human health worldwide, with 4.95 million deaths associated with AMR in 2019 [Murray et al., 2022]. Data on bacterial drug resistance is often either inferred from genome information (leading to uncertainty) or measured using laborious experiments (often leading to incompleteness). The ability to handle incomplete, phylogenetically-related data is therefore a substantial advantage to accumulation models in this context.
@@ 2 Materials and methods
Hypercubic transition paths with Linear Algebra for Uncertainty. We consider a state space $S$ that is the set of all binary strings of length $L$, where $L$ is the number of features in our system. Each binary string describes a state by the presence (1) or absence (0) of each of these $L$ features. A transition from state $s_{i}$ to $s_{j}$ is allowed if the binary string $s_{j}$ differs from $s_{i}$ by the presence of exactly one ‘1’: that is, only transitions that involve the acquisition of one feataure are allowed. The resulting structure is a directed acyclic graph which we call a hypercubic transition graph. The HyperLAU algorithm can be used to learn the transition probabilities $a_{i,j}=\{X_{t}=s_{i}|X_{t-1}=s_{j}\}$ of a discrete Markov process $X_{t}$ on such a hypercubic transition graph.
The likelihood calculation finds the probability that a random walk that runs from the $000$ state to the $111$ state and emits exactly two signals compatible with the ancestral and descendant observations in a datapoint. Beginning with an initial guess of a transition matrix, we then optimize the likelihood.
@@
HyperLAU allows for inference of the transition parameters under different models for interaction or independence between features involved. The user can choose what level of interactions or influences between the different features they want to allow. This is done by an input parameter model$\in\{F,1,2,3,4\}.$ For model $=F$ (full), arbitrary interactions between all features are allowed. In this case, we learn a transition parameter for every allowed state-to-state transition by itself. We consider the whole transition matrix directly as the object we change in order to optimize the likelihood. For the four other models, state-state transition probabilities are functions of the changing feature and a set of contributions from acquired features, as also used in [Aga et al., 2024]. This means our parameter set consists of a base rate with which every feature occurs, and for model$\in\{2,3,4\}$ additional parameters that reflect the influence that the existence of a set of features can have on the rate with which another feature is obtained. In model = $1$, there is only the base rate and it does not matter what other features are already obtained. The other three models allow accordingly interactions of order two, three or four. In these models, influences between either single respectively pairs or triplets of already acquired features can increase or decrease the probability for other features to occur. During the optimization procedure we make changes only on this parameter set. The transition matrix, which still is used in the calculation of the likelihood, can easily be created from this parameter set.
@@
Input: Vectors with ancestor and descendant states of the dataset and the corresponding frequency, length of the strings $L$, transition matrix $A$
@@
[…]
@@
[…]
@@
For a given transition matrix $A$ of the hypercube, we can compute a vector $P_{.,t}$ for every evolutionary step $t=0,...,L$, where $L$ is the length of the strings, i.e. the number of considered features. $P_{s,t}$ is then the probability of a trajectory being in evolutionary step $t$ at node $s$, where $s$ is the natural number that corresponds to the binary string. That means, every $P_{.,t}$ gives a probability distribution of occupancy for the different nodes of the hypercube at step $t$.
Set $P_{.,0}=[1,0,0,...]$. This reflects our assumption that every trajectory has to start in the state 000…, where none of the features have yet been acquired. Then
@@
$Q_{.,t^{\prime}+1}^{t,a}=A\cdot Q_{.,t^{\prime}}^{t,a}\quad\text{ for }t^{\prime}={t,...,L}.$
@@
Calculating Equation (2) for every entry in our dataset, take the logarithm and summing up the obtained log-likelihoods, we get an overall logarithmic likelihood of obtaining the data in the dataset, given the transition matrix. This is also the function that needs to be optimized by adapting the transition or rate matrix.
@@
The initial temperature is set to 1 and decreased in every loop by dividing by a certain factor, which can be specified as an input parameter. For all results here presented, the temperature was divided by 1.001 in every step. As soon as the temperature falls below a threshold of $10^{-6}$, the iterations and with that the optimization process are stopped. This choice of parameter values gave us a sufficient level of convergence for the datasets we considered (for the progression trajectories of the likelihood see SI). All results of the tuberculosis data are obtained under model $F$.
@@
Generating independent observations. Consider the case where we have phylogenetically-embedded observations $0^{L}\rightarrow a,\ a\rightarrow b$ and $a\rightarrow c$, corresponding to observations of $0^{L}$ at the root of the tree, state $a$ at ancestral node $A$, and states $b$ and $c$ at two daughter tips $B$ and $C$. Treating emission probabilities as constants, we can describe the probability of obtaining these observations by:
$P(\text{observe }0^{L}\rightarrow a,a\rightarrow b,a\rightarrow c)P(0^{L}\rightarrow a)P(a\rightarrow b)P(b\rightarrow c)$
@@
$P(\text{observe }0^{L}\rightarrow a,a\rightarrow b,a\rightarrow c)\sum_{a^{\prime}}P(0^{L}\rightarrow a^{\prime})P(a^{\prime}\rightarrow b)P(a^{\prime}\rightarrow c)$
@@
Implementation. The code is implemented in C++ and is freely available on Github: https://github.com/JessicaRenz/HyperLAU. The Armadillo library for linear algebra and scientific computing [Sanderson and Curtin, 2016, Sanderson and Curtin, 2019] is used in this code. Additionally there are also R scripts for creating the visualisations of the output, string manipulation and tree processing, which use the libraries ggplot2 [Wickham, 2016], stringr [Wickham, 2023], ggraph [Pedersen, 2024], ggpubr [Kassambara, 2023], igraph [Csardi and Nepusz, 2006, Csárdi et al., 2024], ggrepel [Slowikowski, 2024], readxl [Wickham and Bryan, 2025], tidyr [Wickham et al., 2024] and dplyr [Wickham et al., 2023].
@@ 3.2 HyperLAU on tuberculosis drug resistance data with artificially introduced uncertainties
Considering the results of the data with 50% inserted uncertainty, the dynamic in the beginning of obtaining resistance first against INH, then against RIF, STR, EMB and PZA is still very clear. However, in this case the version with INH before RIF is clearly favored. It is also visible that several other edges have a higher flux as before, although with a higher CV. This increased uncertainty is to be expected when having more uncertainty. Taken together, despite the challenges involved in artificially obscuring half of the original data, HyperLAU can clearly capture the most likely and outstanding pathways of a real-world data set.
3.3 HyperLAU on tuberculosis drug resistance data with naturally included uncertainty
To demonstrate how HyperLAU’s capacity for uncertain data allows an expansion beyond previous analysis methods, we applied HyperLAU to the full original tuberculosis dataset from [Casali et al., 2014], including samples for which is no information for some features available. These uncertain samples have been excluded in our previous example, and also in the applications of other methods like HyperHMM and HyperTraPS before [Moen and Johnston, 2023, Aga et al., 2024]. With HyperLAU it is now possible to take the full set of 1000 samples into account (see Methods for a discussion of how phylogenetically independent transitions were obtained), and get a more complete evolutionary picture than before.
Figure 4: Inference of anti-microbial evolution in tuberculosis based on a full dataset including uncertainties. (A) Visualisation of the used dataset from [Casali et al., 2014] embedded in a phylogeny. Each row in the matrix corresponds to a bacterial isolate that is a tip in the phylogeny. Each column in the matrix describes resistance to a different drug: red fields in the profile represent missing data, white fields indicate the absence of resistance, black fields indicate the presence of resistance. (B) Illustration of the transitions (not the individual isolate profiles) present in this dataset. For each drug, pie segment colour describes the ‘before’ state of that drug in a transition, and the labels on the circumference describe the ‘after’ state. The size of each pie section gives the proportion of transitions for that before-after combination. (C) Transition network visualisation of the fluxes learned by HyperLAU through the evolutionary state space from 000… (top) to 111… (bottom) (as in Figure 1D). The thickness of the edges represents the strength of the flux. The coefficient of variation (CV) is illustrated by the colour. Edges are labelled by the drug resistance that is gained in that step. Only edges with a flux $>0.01$ are shown. Illustration of the bootstrap uncertainty for the first steps is given in Figure S5.

### v3

@@ Abstract
Understanding and predicting evolutionary accumulation pathways is a key objective in many fields of research, ranging from classical evolutionary biology to diverse applications in medicine. In this context, we are often confronted with the problem that data is sparse and uncertain, which challenges many existing approaches for inferring evolutionary pathways. To use the available data as powerfully as possible, inference approaches that can handle this uncertainty are required. In this article we introduce HyperLAU, a new algorithm for learning evolutionary pathways for accumulation processes from datasets with substantial uncertainty. Expanding the flexibility of accumulation modelling, HyperLAU allows us to infer dynamic pathways and interactions between features, even when large sets of particular features are unobserved across the source dataset. We show that HyperLAU is able to identify evolutionary pathways found by other tools that rely on completely specified data, even when up to 50% of the features in the input data are artifically obscured. Additionally, we demonstrate how it can help to overcome possible biases that can occur when reducing the used data by excluding uncertain subsets. We illustrate the approach with case studies on the evolution of multidrug resistance in tuberculosis, $C_{4}$ photosynthesis in plants, and mitochondrial reduction across eukaryotes, showing in each case that HyperLAU allows more flexible data, strengthens scientific interpretation, and provides new information about evolutionary pathways compared to existing methods.
@@ 1 Introduction
Many processes in the natural sciences and especially in biology and medicine can be modelled as the accumulation or loss of binary features [Diaz-Uriarte and Johnston, 2025, Schill et al., 2024b, Diaz-Uriarte and Herrera-Nieto, 2022]. Common examples include evolutionary processes, such as accumulation of mutations in cancer [Schwartz and Schäffer, 2017, Diaz-Uriarte and Vasallo, 2019, Beerenwinkel et al., 2015, Takeshima and Ushijima, 2019, Moen and Johnston, 2023], gene loss in mitochondria [Johnston and Williams, 2016, Maier et al., 2013], mutations in bacteria [Nichol et al., 2015, Tan et al., 2011, Greenbury et al., 2020, Aga et al., 2025], and the appearance of other phenotypic characters [Moen and Johnston, 2023, Nichol et al., 2015]. Other examples of accumulation processes include, for example, patients acquiring symptoms in progressing diseases [Dalgic et al., 2021, Johnston et al., 2019, Schill et al., 2024a]. Understanding and predicting such processes can contribute both to the basic knowledge of the mechanisms involved, and to potential application – for example, in triaging patients [Johnston et al., 2019] and determining optimal treatments [Aga et al., 2024, Renz et al., 2025].
To address this problem, many different types of mathematical models for accumulation processes have been derived over the last decades, many of them using discrete binary structures. Among the first models in this context, the Mk (Markov k-state) model [Pagel, 1994, Lewis, 2001] for the evolution of traits across a group of related species was introduced. Recently, [Johnston and Diaz-Uriarte, 2024] demonstrated how the Mk model can be used to infer accumulation processes, including the possibility of reversible transitions.
@@
In several scientific and medical contexts, from drug-resistance phenotypes to anatomical characters, data on evolutionary accumulation processes often includes uncertainties and unknown states. The extent to which existing models can handle this uncertainty is limited. In most of the named models, there are only general correction mechanisms or error rates included, which can account for some noise. MO [Gao et al., 2022] explicitly allows missing or ambiguous information for specific features, but is subject to the infinite-sites assumption (every event occurs at most once, prohibiting parallel evolution), which is not appropriate in several scientific contexts. HyperTraPS-CT [Aga et al., 2024] does not have this assumption, but represents data as a set of ancestor-descendant states and can only naturally deal with uncertainties in the descendant states, not in the ancestral ones. So far the only model that can fulfill all these requirements is the Mk-model [Johnston and Diaz-Uriarte, 2024] which, however, is limited to very few features due to the high computational runtime.
Here, we introduce HyperLAU (Hypercubic transition paths with Linear Algebra for Uncertainty), an approach for accumulation modelling that has no underlying infinite site assumption but can deal with longitudinal input data, including specified uncertainties in ancestor and descendant states. On the same time, it can handle input data with more features than the Mk-model in a smaller amount of time. Pursuing the idea of an hypercubic transition graph and maintaining the option of longitudinal and/or phylogenetic data as input, we developed an approach where precise positions of uncertainty can be specified in both the ancestor and descendant strings. In the center of the algorithm are simple matrix-vector multiplications, as known from linear algebra, which makes the approach intuitive and easy understandable (see Figure 1). Borrowing the idea for the option to model different orders of dependencies from [Greenbury et al., 2020], we additionally keep this flexibility, constituting a further advantage of our model. In addition to case studies on diverse evolutionary topics from photosynthesis to organelle evolution, we will particularly motivate the approach in the context of anti-microbial resistance (AMR) evolution – the evolution of pathogens to different drugs. AMR is one of the top threats to human health worldwide, with 4.95 million deaths associated with AMR in 2019 [Murray et al., 2022]. Data on bacterial drug resistance is often either inferred from genome information (leading to uncertainty) or measured using laborious experiments (often leading to incompleteness). The ability to handle incomplete, phylogenetically-related data is therefore a substantial advantage to accumulation models in this context [Renz et al., 2025, Aga et al., 2025].
@@ 2 Materials and methods
Hypercubic transition paths with Linear Algebra for Uncertainty. We consider a state space $S$ that is the set of all binary strings of length $L$, where $L$ is the number of features in our system. Each binary string describes a state by the presence (1) or absence (0) of each of these $L$ features. A transition from state $s_{i}$ to $s_{j}$ is allowed if the binary string $s_{j}$ differs from $s_{i}$ by the presence of exactly one ‘1’: that is, only transitions that involve the acquisition of one feataure are allowed. The resulting structure is a directed acyclic graph which we call a hypercubic transition graph. The HyperLAU algorithm can be used to learn the transition probabilities $a_{i,j}=\{X_{t}=s_{i}|X_{t-1}=s_{j}\}$ of a discrete Markov process $X_{t}$ on such a hypercubic transition graph. In so doing, the likely evolutionary pathways supported by observation are inferred.
The likelihood calculation finds the probability that a random walk that runs from the 000000... state to the 111111... state and emits exactly two signals compatible with the ancestral and descendant observations in a datapoint. Beginning with an initial guess of a transition matrix, we then optimize the likelihood.
@@
[…]
@@
Input: Vectors with ancestor and descendant states of the dataset and the corresponding frequency, length of the strings $L$, transition matrix AA
@@
18:       if sum($Q_{.,t^{\prime}}^{\text{comp}}$) $\neq 0$ then
19:         store sum($Q_{.,t^{\prime}}^{\text{comp}}$)
@@
24:  $\mathcal{L}^{\prime}=$ sum of all stored sum($Q_{.,t^{\prime}}^{\text{comp}}$) multiplied with normalization factor
@@
For a given transition matrix $A$ of the hypercube, we can compute a vector $P_{.,t}$ for every evolutionary step t=0,,Lt=0,...,L, where $L$ is the length of the strings, i.e. the number of considered features. $P_{s,t}$ is then the probability of a trajectory being in evolutionary step $t$ at node $s$, where $s$ is the natural number that corresponds to the binary string. That means every $P_{.,t}$ gives a probability distribution of occupancy for the different nodes of the hypercube at step $t$. First we set $P_{.,0}=[1,0,0,$...]. This reflects our assumption that every trajectory has to start in the state 000…, where none of the features have yet been acquired. Then
@@
$Q_{.,t^{\prime}+1}^{t,a}=A\cdot Q_{.,t^{\prime}}^{t,a}\qquad\text{ for }t^{\prime}={t,...,L}.$
@@
Calculating Equation (2) for every entry in our dataset, taking the logarithm, and summing up the obtained log-likelihoods, we get an overall logarithmic likelihood of obtaining the data in the dataset, given the transition matrix. This is also the function that needs to be optimized by adapting the transition or rate matrix.
@@
The initial temperature is set to 1 and decreased in every loop by dividing by a certain factor, which can be specified as an input parameter. For all results here presented, the temperature was divided by 1.001 in every step. As soon as the temperature falls below a threshold of $10^{-6}$, the iterations and with that the optimization process are stopped. This choice of parameter values gave us a sufficient level of convergence for the datasets we considered (for the progression trajectories of the likelihood see SI). All results of the tuberculosis, photosynthesis, and mitochondrial evolution case studies are obtained under model $F$.
@@
Generating independent observations. Consider the case where we have phylogenetically embedded observations $0^{L}\rightarrow a,\ a\rightarrow b$ and $a\rightarrow c$, corresponding to observations of $0^{L}$ at the root of the tree, state $a$ at ancestral node $A$, and states $b$ and $c$ at two daughter tips $B$ and $C$. Treating emission probabilities as constants, we can describe the probability of obtaining these observations by:
[…]
@@
[…]
@@
Implementation. The code is implemented in C++ and is freely available on Github: https://github.com/JessicaRenz/HyperLAU. The Armadillo library for linear algebra and scientific computing [Sanderson and Curtin, 2016, Sanderson and Curtin, 2019] is used in this code. Additionally there are also R scripts for creating the visualisations of the output, string manipulation and tree processing, which use the libraries ggplot2 [Wickham, 2016], stringr [Wickham, 2023], ggraph [Pedersen, 2024], ggpubr [Kassambara, 2023], igraph [Csardi and Nepusz, 2006, Csárdi et al., 2024], ggrepel [Slowikowski et al., 2018], readxl [Wickham et al., 2019], tidyr [Wickham and Wickham, 2017] and dplyr [Wickham et al., 2023].
@@ 3.2 HyperLAU on tuberculosis drug resistance data with artificially introduced uncertainties
Considering the results of the data with 50% inserted uncertainty, the dynamic in the beginning of obtaining resistance first against INH, then against RIF, STR, EMB and PZA is still very clear. However, in this case the version with INH before RIF is clearly favored. It is also visible that several other edges have a higher flux as before, although with a higher CV. This increased uncertainty in inference output is not unexpected given associated uncertainty in the data input. Taken together, despite the challenges involved in artificially obscuring half of the original data, HyperLAU can clearly capture the most likely and outstanding pathways of a real-world data set.
3.3 Case study 1: Uncertainty in resistance classifications in tuberculosis multi-drug resistance evolution
To demonstrate how HyperLAU’s capacity for uncertain data allows an expansion beyond previous analysis methods, we applied HyperLAU to the full original tuberculosis dataset from [Casali et al., 2014], including samples for which is no information for some features available (due to the computational difficulty of inferring drug resistances from genome data). These uncertain samples have been excluded in our previous example, and also in the applications of other methods like HyperHMM and HyperTraPS before [Moen and Johnston, 2023, Aga et al., 2024]. With HyperLAU it is now possible to take the full set of 1000 samples into account (see Methods for a discussion of how phylogenetically independent transitions were obtained), and get a more complete evolutionary picture than before.
Figure 4: Inference of anti-microbial evolution in tuberculosis based on a full dataset including uncertainties. (A) Visualisation of the used dataset from [Casali et al., 2014] embedded in a phylogeny. Each row in the matrix corresponds to a bacterial isolate that is a tip in the phylogeny. Each column in the matrix describes resistance to a different drug: red fields in the profile represent missing data, white fields indicate the absence of resistance, black fields indicate the presence of resistance. (B) Illustration of the transitions (not the individual isolate profiles) present in this dataset. For each drug, pie segment colour describes the ‘before’ state of that drug in a transition, and the labels on the circumference describe the ‘after’ state. The size of each pie section gives the proportion of transitions for that before-after combination. (C) Transition network visualisation of the fluxes learned by HyperLAU through the evolutionary state space from 000… (top) to 111… (bottom) (as in Figure 1D). The thickness of the edges represents the strength of the flux. The coefficient of variation (CV) is illustrated by the colour. Edges are labelled by the drug resistance that is gained in that step. Only edges with a flux >0.01>0.01 are shown. Illustration of the bootstrap uncertainty for the first steps is given in Figure S5.
@@
3.4 Case study 2: Uncertainty in phenotypic features in $C_{4}$ photosynthesis evolution
An early application of EvAM was the remarkable convergent evolution of $C_{4}$ photosynthesis in plants [Williams et al., 2013]. Involving the acquisition of a collection of physiological, cellular, and gene expression adaptations, $C_{4}$ has evolved independently over 60 times. Previous work used accumulation modelling to develop a ‘roadmap’ for $C_{4}$ evolution across these instances, using data from $C_{3}-C_{4}$ intermediates’ – species that have acquired some but not all of the features involved in $C_{4}$. These results partly explained the remarkably widespread occurrence of $C_{4}$ by identifying that early steps in its evolution were largely coarse-grained physiological ones that plausibly have positive fitness consequences even in the absence of other $C_{4}$ traits, ‘paving the road’ for further refinement [Williams et al., 2013].
Due to the technical difficulties of phenotypic observation in diverse $C_{3}-C_{4}$ intermediates, much (over 59%) of the original dataset was incomplete. At the time, this require a naive sampling approach where walkers were simulated in state space and with no guarantee that they would usefully contribute to likelihood estimation, then recorded if they encountered a state compatible with a given incomplete observation. The uncertainty in the resultant roadmap was substantial, and the computational expense was considerable. HyperLAU’s natural uncertainty embedding allows this to be addressed much more satisfactorily.
Fig. 5A shows the results of HyperLAU inference on the first ten features in this dataset. The first two steps – enlarged bundle sheath cells and cell specificity of GDC – remain overwhelmingly favoured as the first evolutionary steps towards $C_{4}$. Interestingly, the next most likely step (albeit with uncertainty) is predicted to be an increase in the number of bundle sheath chloroplasts. This is a coarse physiological change but was not ranked as an early precursor in the original paper because its ordering posterior was rather bimodal, with density both at early and late acquisition orderings. A similar effect is seen for GDC abundance, here inferred to be earlier than in the original study, and with a bimodal distribution in that study. The explicit handling of uncertainty has here reinforced the scientific message – that coarse physiological changes, of benefit outside the specific $C_{4}$ context, are early steps ‘paving the way’ for full $C_{4}$ evolution.
Figure 5: HyperLAU case studies in photosynthesis and reductive mitochondrial evolution. Transition networks inferred using HyperLAU, styled as in Fig. 4C. A: $C_{4}$ photosynthesis case study, with data from [Williams et al., 2013] (and references therein). Features (where ‘specificity’ means cell-type specificity: (1) large bundle sheath (BS) cells; (2) GDC cell specificity; (3) vein spacing; (4) decarboxylase specificity; (5) PPDK specificity; (6) BS chloroplast number; (7) GDC abundance; (8) PPDK abundance); (9) decarboxylase abundance; (10) RuBisCO abundance. B: Mitochondrial reduction case study, with data from [Glastad and Johnston, 2025] (and references therein). Features: electron transport chain complexes (1) CI; (2) CII; (3) CIII; (4) CIV; (5) CV; (6) pyruvate dehydrogenase; (7) mtDNA; (8) citric acid cycle steps; (9) iron-sulfur metabolism.
3.5 Case study 3: Uncertainty in bioenergetic features in reductive mitochondrial evolution
A recent study has used EvAM to characterise the pathways of reductive mitochondrial evolution [Glastad and Johnston, 2025]. Here, in specific (often parasitic) lineages across eukaryotes, mitochondria have lost one or most aspects of their canonical functionality [Hampl and Roger, 2024, Stairs et al., 2015, Maciszewski and Karnkowska, 2019]. This reduction to mitochondrion-related organelles or MROs may involve the loss of one or more electron transport chain complexes, the mitochondrial DNA, or the very ability to produce ATP (iron-sulfur cluster metabolism seems to be the most conserved feature of MROs). The study in [Glastad and Johnston, 2025] showed that the diversity of MROs across eukaryotes can largely be explained by ‘variations on a theme’ of two distinct pathways: the CI-pathway, involving loss of Complex I before other features; and the alt-pathway, involving loss of either CIII, CIV, or steps of the citric acid cycle before other features.
However, again, the characterisation of the presence and absence of molecular features of mitochondria in diverse eukaryotes (including many protists) is a challenging task. The [Glastad and Johnston, 2025] study surveyed the literature to construct a dataset for the EvAM approach. However, several inconsistencies came up across literature sources, with 26 out of 81 organisms containing some conflicting information about presence/absence of features. In the original study this diversity was condensed into two datasets reflecting the ‘uncertainty means absence’ and ‘uncertainty means presence’ limits, and inferences drawn about results that remained consistent across the two. But HyperLAU offers a much more satisfactory approach of treating conflicting observations as uncertain.
The results are shown in Fig. 5B. The structuring in CI- and alt-pathways is still visible, but the CI-pathway is dramatically favoured compared to the alt-pathway in this HyperLAU treatment. The structure of the alt-pathway is more canalised under HyperLAU, with CIV as the most common initial loss step and a previously unidentified mechanism – loss of CIV, then CI, then following the CI-pathway – showing some consistent probability. The outcome of accounting for uncertainty in this case has been to reweight the previous inference outcomes, providing a more parsimonious explanation of the diverse dynamics observed across eukaryotes: the highly probable CI-pathway is highly canalised and predictable, and tied to limiting metabolic impact of mitochondrial reduction [Glastad and Johnston, 2025].